Find the square root of โˆ’5 โˆ’ 12๐‘–

1
Expert's answer
2022-03-21T00:38:17-0400

"-5-12i=\\left( a+ib \\right) ^2\\\\-5-12i=a^2-b^2+2abi\\\\\\left\\{ \\begin{array}{c}\ta^2-b^2=-5\\\\\t2ab=-12\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tb=-\\frac{6}{a}\\\\\ta^2-\\frac{36}{a^2}=-5\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tb=-\\frac{6}{a}\\\\\ta^4+5a^2-36=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tb=-\\frac{6}{a}\\\\\ta=\\pm 2\\\\\\end{array} \\right. \\Rightarrow \\\\\\Rightarrow \\left\\{ \\begin{array}{c}\t\\sqrt{-5-12i}_1=2-3i\\\\\t\\sqrt{-5-12i}_2=-2+3i\\\\\\end{array} \\right."